# ACPO: AI-Enabled Compiler-Driven Program Optimization

### 15 Dec 2023 v1

@@ Abstract.
The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler. Traditionally, such profitability decisions are made by hand-coded algorithms tuned for a very small number of benchmarks, usually requiring a great deal of effort to be retuned when the benchmark suite changes. This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework. Even so, seamless integration of ML into the compiler would require constant rebuilding of the compiler when models are updated.
This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% and 3%, 5.4%, 0.2% compared to LLVM’s O3 optimization when deployed on Polybench, Coral-2, CoreMark, and Graph-500, respectively. Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
@@
copyright: nonejournal: tacoccs: Software and its engineering Compilersccs: Computing methodologies Machine learning{textblock}
13(1.6,1) Under revision at ACM Transactions on Architecture and Code Optimization (TACO)
Received 26 June 2023; Revised 20 Nov 2023
LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
@@ 1. Introduction
Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years. In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023). Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a). To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a). In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
Recently, there have been a few works providing an end-to-end approach to construct ML-based heuristics for optimizing the decision making process of the transformation passes (Cummins et al., 2017; Das et al., 2020; Rotem and Cummins, 2021; Trofin et al., 2021; Cummins et al., 2021b; Mannarswamy and Das, 2022; Ashouri et al., 2022). Among these, only a subset provides an ML-guided optimization approach (Cummins et al., 2021b; Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022) baked into a single or multiple passes to readily streamline the optimization process inside the pass functionality. The recent advancements in this area beg the question of how scalable, user-friendly, and comprehensive a framework should be, to provide seamless integration with existing LLVM optimization passes, while maintaining a clean separation of concerns between the ML and the LLVM layers.
@@
The rest of the paper is organized as follows. Section 2 discusses the state of the art. Section 3 gives details on our proposed framework and how the different layers of abstraction communicate with each other. Under Section 4.1, we showcase how Loop Unroll pass can benefit by instantiating ACPOLUModel and, subsequently, Section 4.2 does so for Function Inlining pass. Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
@@ 2. Related Work
A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018a) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a). Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
@@
Rotem and Cummins (Rotem and Cummins, 2021) propose a Decision Trees (DS)-based framework to infer branch probabilities for use in providing profile-guided optimizations (PGO). The authors use hand-crafted features for the characterization of code regions and show they are able to reap the benefits of PGO without actually running the program with instrumentation to collect profile data and recompiling. This work is relevant to ours since the approach builds into LLVM and once trained, it can provide PGO certain profiling information without actually running PGO.
Trofin et al. (Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Our previous work, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g. SPEC CPU 2006 and Cbench.
@@ 3. Proposed Methodology
3.1. Autotuner
An autotuner is a software tool that derives decisions for an optimization scenario (minimization or maximization), by intelligently exploring the design space of compiler optimization parameters (Palermo et al., 2005; Ansel et al., 2014; Ashouri et al., 2018a, b). The parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014). This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
@@ 3.2. ACPO Training
To generate data for training purposes we rely on the autotuner. The autotuning process allows us to override compiler heuristics with an iterative search for the optimal optimization decisions for a given application. By performing the search on our benchmark programs and recording each step of the search, we can obtain information about the impact of each decision on program performance, and thus collect sufficient information to be able to train the eventual model with a reasonable chance of success. To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another. For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops. With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks.
The first step in the autotuning flow is to establish a search space over which the tuning occurs. This comprises both the focus on a pass as well as code regions on which the specific transformation can act. In loop unrolling for instance, the applicable code regions would be loops, not single instructions or whole functions. Once the search space is established, the autotuner simply proposes solutions for each code region and then requests the compiler to apply the desired optimization on the given code region at the appropriate time. This is crucial because we do not want to affect the compiler pipeline itself; instead, we want to generate a wide array of solutions at a specific point in the compilation process and then observe the impact of those decisions. That is key to creating a viable model.
@@
At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions. Initially, the pass proceeds normally, starting with legality checks and then following through to profitability analysis, which is when the ML model will be invoked. The compiler instantiates an ACPOModel and gathers features pertaining to the code region being optimized. This is important to ensure that the model being trained is provided with input features based on what they will be at deployment time when the compiler uses this ML-enabled pass. Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation. The features are passed with std::vector of pairs to the ML side. This process can also be scaled to an arbitrary number of features and feature types. The feature set is then saved for later use to allow the features, the decision specified by the autotuner, and the final impact of the decision to be correlated. The transformation decision as prescribed by the autotuner is applied to the code region (e.g. the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated.
Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program). The measurement is used as an assessment of decision impact, so the decisions that lead to a positive impact can be distinguished from the ones that lead to a negative impact on performance. Generally, we use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation. This process is repeated many times to gather as much data per benchmark as possible to achieve beneficial results on most benchmarks. The resulting trial data is then processed and used as input to the model training.
@@ 3.3. ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged.
The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016). Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop. To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
@@ 3.4. Compiler
One of the key things we added to the compiler is the ability to invoke a model and run inference given a set of features. To facilitate this, ML-enabled passes in the compiler have been given access to ACPOModel class, from which a pass-specific model instance can be derived.
@@ 3.5. ACPOModel
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference. The header file snippets of ACPOModel can be found in the Appendix section under Listing 7.
@@
The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method. The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run. If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr. This will be a crucial check to verify that ML inference is completed and should be checked by the caller.
It is worth noting that the prepareModelInput() and runModel() methods are virtual. The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
3.6. ML APIs
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler. The snippet for the ACPOMLInterface header is shown under the Appendix section in Listing 8.
@@ 3.6. ML APIs
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). The ML framework in our scenario is a separate python process that has a user-visible interface through the named pipes. It receives commands from ACPOMLInterface-derived class on the compiler side, processes the requests, and responds with requested information or status. This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes. We envision this will grow to comprise a toolkit with many different models, algorithms, and program representation mechanisms at the developer’s disposal. Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 3.
3.7. ML Framework
In this work, to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment. Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects.
@@ 4.1. ACPO Loop Unrolling Model
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018a). There have been multiple works trying to automatically construct heuristics to find the optimal unrolling factor (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018) and research has shown identifying the optimal unroll count is a fine line between the increase in speed versus the binary code size by which the increase in cache misses can be counterproductive. Therefore, a number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.

### 11 Mar 2024 v2

@@ Abstract.
The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler. This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework.
This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% compared to LLVM’s O3 optimization when deployed on Polybench. Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
@@
copyright: noneccs: Software and its engineering Compilersccs: Computing methodologies Machine learning{textblock}
30(2,0.5) LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
@@ 1. Introduction
Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years. In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023) and lately, using Large Language Models (LLMs) (Cummins et al., 2023; Gong et al., 2024). Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a). To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a). In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
@@
The rest of the paper is organized as follows. Section 2 discusses the state of the art. Section 3 gives details on our proposed framework. Under Sections 4.1 and 4.2, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel. Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
@@ 2. Related Work
A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a). Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
@@
Trofin et al. (Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Later, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g. SPEC CPU 2006 and Cbench.
@@ 3. Proposed Methodology
Autotuner
The autotuner parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014). This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
@@ 3.1. ACPO Training
To generate data for training purposes we rely on the autotuner. To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another. For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops. With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks. At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation. The features are passed with std::vector of pairs to the ML side. This process can also be scaled to an arbitrary number of features and feature types. The transformation decision as prescribed by the autotuner is applied to the code region (e.g. the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated. Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program). We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
@@ 3.2. ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged. The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016). Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop. To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
@@ 3.4. ACPOModel
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference. The header file code snippet is left out of this version for brevity, but it contains the same functionality described in figures 1 and 2.
@@
The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method. The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run. If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr. This will be a crucial check to verify that ML inference is completed and should be checked by the caller. It is worth noting that the prepareModelInput() and runModel() methods are virtual. The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
3.5. ML APIs & Framework
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler. The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
@@ 3.5. ML APIs & Framework
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes. Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 1. Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects and in this work to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
@@ 4.1. ACPO Loop Unrolling Model
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018). There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018). A number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.